OpenStudy (anonymous):

Simplify : 2^6x/2/6^2x/5^2

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

Eeek! Parenthesis please... What you literally just wrote was this: \[2^6\frac{\frac{x}{2}}{6^2}\frac{x}{5^2} \] which I don't think is what you intended... right? Did you mean this? \[\huge \frac{\frac{2^6x}{2}}{\frac{6^2x}{5^2}} \]

OpenStudy (saranya):

sry the ans is 22.22

OpenStudy (anonymous):

\frac{numerator}{denominator} \[\frac{numerator}{denominator}\] \huge \frac{\frac{2^6x}{2}}{\frac{6^2x}{5^2}} ((2^6x)/(2))/((6^2x)/(5^2))

OpenStudy (saranya):

2^6x÷2×5^2÷6^2x=2^5×5^2÷6^2=22.22

OpenStudy (anonymous):

@saranya you're making an assumption as to what was the intended problem, and the variable x doesn't just disappear...

OpenStudy (saranya):

is t not 2^6 * x / 2

OpenStudy (saranya):

No variable x gets cancelled !

OpenStudy (anonymous):

* or \(\cdot\) is multiplication x is a letter variable, typically, sometimes can be mistaken for multiplication I might get canceled or it might not. Given what you assumed, I would agree with you, yes. ;-) But I can read both what the author of the question literal wrote versus what they probably intended to write if they forgot a few parenthesis to make the order of operations go as intended. @XBiyX Was that literal?

OpenStudy (saranya):

yea d same thing......... i meant to say 2^6 x / 2 = (64x)/2